Ballery1:

Find the derivative by using the basic rules of differentiation F(x)=√2

Ballery1:

Find the derivative by using the basic rules of differentiation \[F(x) = \sqrt{2}\]

Ballery1:

Let me draw my solution plz

Ballery1:



Ballery1:

@Hero

katwhirles:

it looks good to me

Ballery1:

Except its not....according to my textbook

katwhirles:

then whats wrong with it

katwhirles:

bc it looks good

Ballery1:

The answer is zero

Vocaloid:

Is the function just f(x) = sqrt(2)? Try writing it as f(x) = sqrt(2) * x^(0) Bring the constant sqrt(2) to the front and apply the differentiation rule to x^(0)

Ballery1:

Yes

Ballery1:

Thanks for ur help voca :)

Ballery1:

I got it. Thanks

Vocaloid:

For future reference all constants have derivative 0 Since a derivative represents change over time, it makes sense for a constant to have 0 change

Ballery1:

Yes i see now. Thanks 🙏🏻

Narad:

If you function is \[f(x)=\sqrt{2}\] Then, you are differentiating a constant.

Ballery1:

Yes that’s the function.

Narad:

ok, then you know the derivative of a constant.

Ballery1:

0

Narad:

ok

Ballery1:

Thanks buddy. Really appreciated your help :)